# The Epistemic Support-Point Filter: Jaynesian Maximum Entropy Meets Popperian Falsification A Possibilistic Minimax-Entropy Optimality Proof

### 10 Mar 2026 v1

@@ Abstract
The Epistemic Support-Point Filter (ESPF) was designed around a single epistemological commitment: be quick to embrace ignorance and slow to assert certainty. This paper proves that this commitment has a precise mathematical form and that the ESPF is the unique optimal filter implementing it within the class of epistemically admissible evidence-only filters.
The ESPF operates as the synthesis of two complementary epistemological principles acting at different phases of the recursion. In its propagation phase it enacts Jaynesian maximum entropy [15]: the support spreads as widely as the dynamics and process noise allow, assuming maximal ignorance consistent with the known dynamical constraints. In its measurement-update phase it enacts Popperian falsification [16]: hypotheses are eliminated if and only if the evidence rules them out, using evidence compatibility alone. Any selection rule that incorporates prior possibility is strictly suboptimal and risks systematic race-to-bottom bias over multiple steps. The filter alternates between these two operations at every step: expand maximally, then contract minimally.
The optimality criterion is possibilistic minimax entropy: among all evidence-only survivor selection rules, the ESPF’s minimum-$q$ selection minimizes the maximum possibilistic entropy consistent with the surviving hypotheses—equivalently, it minimizes $\log\det(\mathrm{MVEE})$, the log-volume of the outermost $\alpha$-cut.
Three lemmas establish the result. The Possibilistic Entropy Lemma identifies the possibilistic entropy $H_{\pi}$—the integral of $\log V_{\alpha}$ over $\alpha\in(0,1]$—as the correct ignorance functional and decomposes it into a minimax term and a gradient term. The Possibilistic Cramér–Rao Lemma bounds how rapidly any admissible filter may reduce $H_{\pi}$ per measurement. The Evidence-Optimality Lemma proves that evidence-only selection minimizes $\log\det(\mathrm{MVEE})$ at every $\alpha$-level simultaneously.
Numerical validation over a 2-day, 877-step Level-3 Smolyak run ($M=106$ support points, $n=7$) under both nominal LEO tracking and a stress case (10 m/s crosstrack maneuver ++ 20 m Arecibo range bias) confirms the theorem’s regime structure. The Epistemic Width Monitor (EWM) diagnostic suite reveals that possibilistic stress under sustained model mismatch manifests primarily through necessity saturation and surprisal escalation rather than MVEE sign change, establishing these as the more sensitive early-warning indicators of model–reality divergence. The Kalman filter is recovered in the Gaussian limit.
@@ 1.1 Two Principles, One Filter
The ESPF differs from Bayesian filters in a foundational way: it minimizes worst-case epistemic ignorance rather than expected uncertainty. A Bayesian filter integrates over a prior to produce a posterior; the ESPF selects survivors by evidence alone and minimizes the maximum residual ignorance consistent with those survivors. Where prior information is unreliable or unavailable, this distinction is not stylistic—it is the difference between an epistemically defensible recursion and one that risks compounding prior errors.
@@ 2.1 Possibility Distributions, $\alpha$-Cuts, and Possibilistic Entropy
A possibility distribution over a finite support $\{\chi^{(i)}\}_{i=1}^{M}\subset\mathbb{R}^{n}$ is a function $\pi:\{1,\ldots,M\}\to(0,1]$ with $\max_{i}\pi^{(i)}=1$. The $\alpha$-cut at level $\alpha\in(0,1]$ is
@@
The $\alpha$-cuts are nested: $C_{\alpha^{\prime}}\subseteq C_{\alpha}$ for $\alpha^{\prime}>\alpha$. The volume of the $\alpha$-cut is $V_{\alpha}=c_{n}(\det\Pi_{\alpha})^{1/2}$, where $\Pi_{\alpha}\succ 0$ is the MVEE shape matrix of $\{\chi^{(i)}:i\in C_{\alpha}\}$ and $c_{n}=\pi^{n/2}/\Gamma(n/2+1)$. When $|C_{\alpha}|<n+1$, the MVEE is degenerate; we set $V_{\alpha}=0$ by convention.
@@
The possibilistic entropy of a possibility distribution $\pi$ over $\{\chi^{(i)}\}_{i=1}^{M}$ is
$H_{\pi}=\int_{0}^{1}\log V_{\alpha}\,d\alpha=\int_{0}^{1}\!\left(\tfrac{1}{2}\log\det\Pi_{\alpha}+\kappa_{n}\right)d\alpha,$ (2)
@@
The possibilistic entropy (2) is the geometric analog of the U-uncertainty measure introduced by Higashi and Klir [7] for discrete possibility distributions on ordered sets. Here the ordering is geometric: the $\alpha$-cuts are nested convex bodies in $\mathbb{R}^{n}$, and the integral of log-cardinality over $\alpha$ is replaced by the integral of log-volume over $\alpha$.
@@
$\hat{\pi}^{(i)}_{k}=\min\!\left(\pi^{(i)}_{k|k-1},\,\mathrm{Comp}^{(i)}_{k}\right),\qquad\mathrm{Comp}^{(i)}_{k}=\exp\!\left(-\tfrac{1}{2}q^{(i)}_{k}\right)$, (3)
@@
After the first update, the ESPF’s possibility distribution is non-uniform in general: the anchor hypothesis (most compatible with $y_{k}$) achieves possibility 1, while other survivors achieve strictly smaller values. Uniformity is a transient condition at initialization, not a maintained property. The correct entropy functional for the ESPF at step $k\geq 1$ is $H_{\pi}$ as defined in (2).
@@ 2.2 Epistemically Admissible Filters
    Possibilistic representation. Uncertainty is represented as a normalized possibility distribution $\pi_{k}$ over a finite support $\{\chi^{(i)}_{k}\}_{i=1}^{M}\subset\mathbb{R}^{n}$ with $\max_{i}\pi^{(i)}_{k}=1$.
@@
    Admissibility of the commitment point. The state estimate $\hat{x}_{k}\in\{\chi^{(i)}_{k}\}$.
@@ 2.3 Innovation Geometry and Possibilistic Information Content
$q^{(i)}_{k}=\left\|L_{e}^{-1}(y_{k}-h(\chi^{(i)}_{k|k-1}))\right\|^{2}_{2},\qquad\Pi_{e}=L_{e}L_{e}^{\top}.$ (5)
@@
The aggregate epistemic surprisal at step $k$ is
$\bar{S}_{k}=\operatorname{median}_{i=1,\ldots,M}\tfrac{1}{2}q^{(i)}_{k}$. (6)
The median is used rather than the mean for robustness: during stress events a small fraction of hypotheses may accumulate very large $q^{(i)}_{k}$; the mean would overstate the information content and produce an inflated PCRB exactly when robustness matters most. The possibilistic information content of $y_{k}$ is
@@
$I_{k}=0$ when the evidence is fully consistent with the predicted support; $I_{k}\to 1$ when evidence is maximally falsifying.
@@ 3.1 Lemma 1: Possibilistic Entropy as Ignorance Functional
Let $\pi$ be a normalized possibility distribution over $\{\chi^{(i)}\}_{i=1}^{M}\subset\mathbb{R}^{n}$ with $\pi^{(i)}\in(0,1]$ and $\max_{i}\pi^{(i)}=1$. The possibilistic entropy $H_{\pi}$ defined in (2) satisfies:
@@
    $H_{\pi}=\underbrace{\log V+\kappa_{n}}_{\text{support entropy}}+\underbrace{\int_{0}^{1}\!\log\frac{V_{\alpha}}{V}\,d\alpha}_{\text{gradient entropy}\;\leq\,0}.$ (8)
@@
    Extremes. $H_{\pi}$ is maximized when $\pi$ is uniform ($H_{\pi}=\log V+\kappa_{n}$) and approaches -$\infty$ as possibility collapses to a single hypothesis. The VFI condition (Definition 2.7(iii)) prevents collapse and keeps $H_{\pi}$ finite within $\mathcal{F}$.
@@
(iii) Write $\log V_{\alpha}=\log V+\log(V_{\alpha}/V)$. Since $C_{\alpha}\subseteq\{1,\ldots,M\}$ for all $\alpha$, we have $V_{\alpha}\leq V$ and $\log(V_{\alpha}/V)\leq 0$. Integrating over $\alpha\in[0,1]$ gives (8).
(iv) Collapse to a singleton gives $|C_{\alpha}|=1<n+1$ for all $\alpha>0$, so $V_{\alpha}=0$ by Definition 2.1 and $H_{\pi}\to$-$\infty$. The upper bound follows from (i). ∎
@@ 3.2 Lemma 2: Possibilistic Cramér–Rao Bound
Let $F\in\mathcal{F}$ be any epistemically admissible filter. At step $k$, let $H_{\pi,k|k-1}$ and $H_{\pi,k|k}$ denote the pre- and post-update possibilistic entropies, and $I_{k}$ the possibilistic information content (7). Then
@@
Step 1: Conjunctive update implies nested post-update $\alpha$-cuts. By Definition 2.7(ii), $\hat{\pi}^{(i)}_{k|k}\leq\pi^{(i)}_{k|k-1}$ for all $i$, so $C^{\mathrm{post}}_{\alpha}\subseteq C^{\mathrm{pre}}_{\alpha}$ for all $\alpha$ and $V^{\mathrm{post}}_{\alpha}\leq V^{\mathrm{pre}}_{\alpha}$.
Step 2: Information content bounds the admissible cardinality fraction. The median compatibility across hypotheses is $\widehat{\mathrm{Comp}}_{k}=\exp(-\bar{S}_{k})=1-I_{k}$. By definition of the median, at least half the hypotheses satisfy $\mathrm{Comp}^{(i)}_{k}\geq 1-I_{k}$, i.e., their post-update possibility (bounded below by $\mathrm{Comp}^{(i)}_{k}$ through the min assignment) is at least $1-I_{k}$. Hence the $\alpha$-cut $C^{\mathrm{post}}_{\alpha}$ contains at least $M/2$ hypotheses for all $\alpha\leq 1-I_{k}$.
Step 3: Cardinality fraction to volume fraction. For $\alpha\leq 1-I_{k}$, the post-update $\alpha$-cut retains at least $M/2$ of the $M$ pre-update hypotheses. By the John–Fritz theorem [9] applied to the MVEE, the volume of the MVEE of any subset of $k$ points drawn from a set whose MVEE has volume $V$ satisfies $V_{\mathrm{sub}}\geq(k/M)^{n/2}V$ asymptotically for uniform distributions over the MVEE. Applying this with $k\geq M/2$ and replacing the proportionality constant conservatively by the compatibility threshold gives
$V^{\mathrm{post}}_{\alpha}\geq(1-I_{k})\cdot V^{\mathrm{pre}}_{\alpha}\quad\text{for all }\alpha\leq 1-I_{k}.$ (10)
@@
For the first integral, use (10):
$\int_{0}^{1-I_{k}}\!\log V^{\mathrm{post}}_{\alpha}\,d\alpha\geq\int_{0}^{1-I_{k}}\!\left(\log(1-I_{k})+\log V^{\mathrm{pre}}_{\alpha}\right)d\alpha.$
For the second integral, note $V^{\mathrm{post}}_{\alpha}\geq 0$ and the interval has measure $I_{k}$; in the falsification regime ($\log\det(\mathrm{MVEE})<0$) the $\alpha$-cut volumes are finite and this term contributes at most a bounded, negative quantity that is absorbed by the $n/2$ scaling factor when passing from the scalar bound to the $n$-dimensional volume scaling $(1-I_{k})\mapsto(1-I_{k})^{n/2}$ on determinants.
Combining both terms:
$H_{\pi,k|k}\geq\log(1-I_{k})\cdot(1-I_{k})+\int_{0}^{1}\!\log V^{\mathrm{pre}}_{\alpha}\,d\alpha$
and, applying the $n$-dimensional scaling $\log V_{\alpha}=\frac{1}{2}\log\det\Pi_{\alpha}+\kappa_{n}$ which introduces a factor of $n/2$ when translating a scalar compatibility bound to the determinant inequality $\det\Pi^{\mathrm{post}}_{\alpha}\geq(1-I_{k})^{n}\det\Pi^{\mathrm{pre}}_{\alpha}$, yields
$H_{\pi,k|k}\geq H_{\pi,k|k-1}+\tfrac{n}{2}\log(1-I_{k}),$ (11)
which is (9). ∎
@@
The PCRB (9) bounds the integral of $\log V_{\alpha}$ over all levels simultaneously, not a single $\alpha$-cut. An admissible filter may reduce $V_{\alpha}$ sharply at some levels while barely changing others, but the net reduction in $H_{\pi}$ cannot exceed $\frac{n}{2}\log(1-I_{k})^{-1}$. This is stronger than any pointwise volume bound.
@@
The volume bound (10) applies when the surviving hypotheses are drawn approximately uniformly over the pre-update MVEE, which holds to first order under the Smolyak grid structure. For general, non-uniform possibility distributions the bound holds with $(1-I_{k})^{n/2}$ rather than $(1-I_{k})$ on the right-hand side of (10) due to the John ellipsoid theorem; the final entropy bound (9) is unchanged.

### 12 Mar 2026 v2

@@ Abstract
This paper proves that the Epistemic Support-Point Filter (ESPF) is the unique optimal recursive estimator within the class of epistemically admissible evidence-only filters. The optimality criterion is possibilistic minimax entropy: among all evidence-only survivor selection rules, the ESPF minimizes worst-case integrated ignorance per measurement step, and no other admissible filter can do better. The proof is constructive: it identifies the precise mathematical form of the epistemological commitment — be quick to embrace ignorance, be slow to assert certainty — and shows that the ESPF is its unique implementation. Where Bayesian filters minimize mean squared error and are driven toward an assumed truth, the ESPF minimizes maximum entropy and surfaces what has not been proven impossible — a fundamentally different epistemic commitment with fundamentally different failure modes.
Two results locate this theorem within the broader landscape of estimation theory.
The first is a unification. The possibilistic entropy $H_{\pi}=\int_{0}^{1}\log V_{\alpha}\,d\alpha$ — the ESPF’s optimality criterion — is the log-geometric mean of the $\alpha$-cut volume family $\{V_{\alpha}\}$, i.e., $\log M_{0}(\{V_{\alpha}\})$ in the Hölder mean hierarchy. The Popperian minimax bound $\log\det(\mathrm{MVEE})$ is $\log M_{+\infty}$; the Kalman MMSE criterion is $\log M_{0}$ evaluated under Gaussian $\alpha$-cut geometry. Possibility theory and probability theory are not competing frameworks: they are evaluations of the same ignorance functional $\log M_{0}(\{V_{\alpha}\})$ under different geometries of $\alpha$-cuts. The Kalman filter is the Gaussian specialization of the ESPF’s optimality criterion, not a separate invention.
The second is a diagnostic. Numerical validation over a 2-day, 877-step Level-3 Smolyak run ($M=106$ support points, $n=7$) under nominal LEO tracking and a sustained stress case (10 m/s crosstrack maneuver combined with a 20 m range bias) reveals that possibilistic stress manifests primarily through necessity saturation and surprisal escalation — not through MVEE sign change. The MVEE regime flag, operating at $p=+\infty$ in the Hölder hierarchy, is the coarsest instrument in the Epistemic Width Monitor (EWM) suite. Necessity and surprisal, operating closer to $p=0$, provide earlier and more sensitive warning of model–reality divergence. This is not an empirical observation: it is a direct consequence of the Hölder ordering $M_{0}\leq M_{+\infty}$.
The theorem rests on three lemmas. The Possibilistic Entropy Lemma decomposes $H_{\pi}$ into a minimax term (minimized by Popperian survivor selection) and a gradient term (minimized by compatibility-based possibility assignment). The Possibilistic Cramér–Rao Bound (PCRB) establishes the maximum rate at which any admissible filter may reduce $H_{\pi}$ per measurement. The Evidence-Optimality Lemma proves that minimum-$q$ selection is the unique minimizer of $H_{\pi}$ within the evidence-only class, and that any rule incorporating prior possibility risks race-to-bottom bias and is strictly suboptimal.
@@ 1.1 Two Principles, One Filter
The ESPF differs from Bayesian filters not merely in criterion but in purpose. A Bayesian filter is truth-seeking: it assumes a generative model, integrates over a prior, and produces a posterior whose objective is convergence to the true state in expectation. The prior has a vote at every step. When the prior is wrong, the filter is driven toward a false truth — and may never recover, because each update reinforces the prior’s influence on the next. The ESPF is not truth-seeking. It is impossibility-eliminating. It makes no claim about where truth is. It removes only what the evidence has ruled out; what remains is not an estimate of truth but the set of everything that has not yet been proven impossible. A wrong prior assignment in the ESPF cannot bias the filter toward a false truth — it can only slow the elimination of inadmissible hypotheses, which manifests immediately as excess epistemic width in the EWM. This is the difference between an epistemically defensible recursion and one that risks compounding prior errors.
@@ 2.1 Possibility Distributions, $\alpha$-Cuts, and Possibilistic Entropy
A possibility distribution over a finite support $\{\chi^{(i)}\}^{M}_{i=1}\subset\mathbb{R}^{n}$ is a function $\pi:\{1,\ldots,M\}\to(0,1]$ with $\max_{i}\pi^{(i)}=1$. The $\alpha$-cut at level $\alpha\in(0,1]$ is
@@
The $\alpha$-cuts are nested: $C_{\alpha^{\prime}}\subseteq C_{\alpha}$ for $\alpha^{\prime}>\alpha$. The volume of the $\alpha$-cut is $V_{\alpha}=c_{n}(\det\Pi_{\alpha})^{1/2}$, where $\Pi_{\alpha}\succ 0$ is the MVEE shape matrix of $\{\chi^{(i)}:i\in C_{\alpha}\}$ and $c_{n}=\pi^{n/2}/\Gamma(n/2+1)$. When $|C_{\alpha}|<2n+1$, the MVEE is degenerate; we set $V_{\alpha}=0$ by convention. The threshold $2n+1$ reflects the requirement that the point cloud span $\mathbb{R}^{n}$ with antipodal coverage in each coordinate direction—the minimum structure for a full-rank MVEE shape matrix $\Pi_{\alpha}$. This is consistent with the filter’s minimum survivor count $N_{\min}=2n+1$ (Section 7), making $N_{\min}$ a direct consequence of $H_{\pi}$’s well-posedness rather than an independent design choice.
@@
The possibilistic entropy of a possibility distribution $\pi$ over $\{\chi^{(i)}\}^{M}_{i=1}$ is
$H_{\pi}=\int_{0}^{1}\log V_{\alpha}\,d\alpha=\int_{0}^{1}\left(\frac{1}{2}\log\det\Pi_{\alpha}+\kappa_{n}\right)d\alpha,$ (2)
@@
The possibilistic entropy is the geometric analog of the U-uncertainty measure introduced by Higashi and Klir [7] for discrete possibility distributions on ordered sets. Here the ordering is geometric: the $\alpha$-cuts are nested convex bodies in $\mathbb{R}^{n}$, and the integral of log-cardinality over $\alpha$ is replaced by the integral of log-volume over $\alpha$.
@@
$\hat{\pi}^{(i)}_{k}=\min\!\left(\pi^{(i)}_{k|k-1},\,\mathrm{Comp}^{(i)}_{k}\right),\quad\mathrm{Comp}^{(i)}_{k}=\exp\!\left(-\tfrac{1}{2}q^{(i)}_{k}\right)$, (3)
@@
After the first update, the ESPF’s possibility distribution is non-uniform in general: the anchor hypothesis (most compatible with $y_{k}$) achieves possibility 1, while other survivors achieve strictly smaller values. Uniformity is a transient condition at initialization, not a maintained property. The correct entropy functional for the ESPF at step $k\geq 1$ is $H_{\pi}$ as defined above.
@@ 2.2 Epistemically Admissible Filters
    Possibilistic representation. Uncertainty is represented as a normalized possibility distribution $\pi_{k}$ over a finite support $\{\chi^{(i)}_{k}\}^{M}_{i=1}\subset\mathbb{R}^{n}$ with $\max_{i}\pi^{(i)}_{k}=1$.
@@
    Admissibility of the commitment point. The operational output $\hat{x}_{k}\in\{\chi^{(i)}_{k}\}$. This is not a claim that $\hat{x}_{k}$ is truth: it is the least inadmissible hypothesis — the support point most compatible with the accumulated evidence. The ESPF makes no assertion about where truth is; it only identifies what has not yet been ruled out, and among those, which is most consistent with the evidence.
@@
Remark 2.8 (Why evidence-referencing is the right admissibility condition).
Condition (v) is not a convenient restriction — it is the epistemological content of the class. A filter that incorporates prior possibility into its survivor selection ranking is not merely using more information; it is making a different kind of claim. It is asserting that hypotheses which were previously plausible deserve preferential treatment over hypotheses that are more consistent with the current evidence. This is the truth-seeking posture: the prior encodes a belief about where truth is, and that belief is allowed to protect hypotheses from elimination even when the evidence argues against them.
The race-to-bottom failure mode (Lemma 3.7, Step 5) is the formal consequence: over multiple steps, prior-based selection concentrates the possibility distribution on hypotheses that are systematically less evidence-consistent. The filter converges — but toward the prior’s truth, not the evidence’s truth. In an impossibility-eliminating filter, this failure mode is not just suboptimal; it is a category error. The prior has no vote on what is impossible. Only the evidence does. Condition (v) enforces this.
@@ 2.3 Innovation Geometry and Possibilistic Information Content
$q^{(i)}_{k}=\left\|L_{e}^{-1}\bigl(y_{k}-h(\chi^{(i)}_{k|k-1})\bigr)\right\|_{2}^{2},\quad\Pi_{e}=L_{e}L_{e}^{\top}.$ (5)
@@
The aggregate epistemic surprisal at step $k$ is the Choquet integral of the per-hypothesis surprisal $\frac{1}{2}q^{(i)}_{k}$ with respect to the prior possibility distribution $\pi_{k|k-1}$ as capacity:
$\bar{S}_{k}=(C)\int\tfrac{1}{2}q^{(i)}_{k}\,d\pi_{k|k-1}=\sup_{i=1,\ldots,M}\min\!\left(\tfrac{1}{2}q^{(i)}_{k},\,\pi^{(i)}_{k|k-1}\right)$. (6)
The second equality is the standard reduction of the Choquet integral under a possibility measure [5]. The capacity is the prior possibility $\pi_{k|k-1}$, not the current compatibility: this keeps the information content epistemically grounded in accumulated evidence history rather than the current measurement alone, and ensures that hypotheses with low prior credibility cannot inflate $\bar{S}_{k}$ through large but epistemically cheap surprisal. The survivor selection rule (Lemma 3.7) remains pure evidence-only minimum-$q$; the prior possibility enters only here, in the bounding quantity.
The possibilistic information content of $y_{k}$ is
@@
$I_{k}=0$ when the evidence is fully consistent with the predicted support; $I_{k}\to 1$ when the most epistemically credible hypothesis is maximally surprised by the measurement.
Remark 2.11 (Choquet integral under a possibility measure).
For a possibility measure $\Pi$ (max-decomposable, $\Pi(A\cup B)=\max(\Pi(A),\Pi(B))$), the Choquet integral of a non-negative function $f$ reduces to $(C)\int f\,d\Pi=\sup_{i}\min(f(i),\pi(i))$ [5]. This is the possibilistic expectation in the sense of Dubois and Prade: it asks for the highest level $t$ at which there exists a hypothesis that is both highly credible ($\pi^{(i)}\geq t$) and genuinely surprised ($f(i)\geq t$). It is robust to outliers by a different mechanism than the median: low-prior-possibility hypotheses are downweighted by the capacity function rather than by a rank cutoff, which is the epistemically correct weighting in a possibilistic framework.
@@ 3.1 Lemma 1: Possibilistic Entropy as Ignorance Functional
Let $\pi$ be a normalized possibility distribution over $\{\chi^{(i)}\}^{M}_{i=1}\subset\mathbb{R}^{n}$ with $\pi^{(i)}\in(0,1]$ and $\max_{i}\pi^{(i)}=1$. The possibilistic entropy $H_{\pi}$ satisfies:
@@
    $H_{\pi}=\underbrace{\log V+\kappa_{n}}_{\text{support entropy}}+\underbrace{\int_{0}^{1}\log\frac{V_{\alpha}}{V}\,d\alpha}_{\text{gradient entropy}\leq 0}.$ (8)
@@
    Extremes. $H_{\pi}$ is maximized when $\pi$ is uniform ($H_{\pi}=\log V+\kappa_{n}$) and approaches -$\infty$ as possibility collapses to a single hypothesis. The VFI condition prevents collapse and keeps $H_{\pi}$ finite within $\mathcal{F}$.
@@
(iii) Write $\log V_{\alpha}=\log V+\log(V_{\alpha}/V)$. Since $C_{\alpha}\subseteq\{1,\ldots,M\}$ for all $\alpha$, we have $V_{\alpha}\leq V$ and $\log(V_{\alpha}/V)\leq 0$. Integrating over $\alpha\in[0,1]$ gives the decomposition.
(iv) Collapse to a singleton gives $|C_{\alpha}|=1<2n+1$ for all $\alpha>0$, so $V_{\alpha}=0$ by convention and $H_{\pi}\to$-$\infty$. The upper bound follows from (i). ∎
@@ 3.2 Lemma 2: Possibilistic Cramér–Rao Bound
Let $F\in\mathcal{F}$ be any epistemically admissible filter. At step $k$, let $H_{\pi,k|k-1}$ and $H_{\pi,k|k}$ denote the pre- and post-update possibilistic entropies, and $I_{k}$ the possibilistic information content. Then
@@
Step 1: Conjunctive update implies nested post-update $\alpha$-cuts. By the admissibility definition, $\hat{\pi}^{(i)}_{k|k}\leq\pi^{(i)}_{k|k-1}$ for all $i$, so $C^{\mathrm{post}}_{\alpha}\subseteq C^{\mathrm{pre}}_{\alpha}$ for all $\alpha$ and $V^{\mathrm{post}}_{\alpha}\leq V^{\mathrm{pre}}_{\alpha}$.
Step 2: Choquet information content bounds the possibility measure of the surviving set. Let $i^{*}$ be the hypothesis achieving the supremum in $\bar{S}_{k}$, so that
$\min\!\left(\tfrac{1}{2}q^{(i^{*})}_{k},\,\pi^{(i^{*})}_{k|k-1}\right)=\bar{S}_{k}.$
This means $\pi^{(i^{*})}_{k|k-1}\geq\bar{S}_{k}$ and $\mathrm{Comp}^{(i^{*})}_{k}=e^{-\frac{1}{2}q^{(i^{*})}_{k}}\geq e^{-\bar{S}_{k}}=1-I_{k}$. Under the conjunctive update, the post-update possibility of $i^{*}$ satisfies
$\hat{\pi}^{(i^{*})}_{k|k}=\min\!\left(\pi^{(i^{*})}_{k|k-1},\,\mathrm{Comp}^{(i^{*})}_{k}\right)\geq\min(\bar{S}_{k},1-I_{k})>0.$
Hence $i^{*}\in C^{\mathrm{post}}_{\alpha}$ for all $\alpha\leq\min(\bar{S}_{k},1-I_{k})$, and the possibility measure of the surviving set satisfies $\Pi_{k|k-1}(C^{\mathrm{post}}_{\alpha})\geq\pi^{(i^{*})}_{k|k-1}\geq\bar{S}_{k}$ for all $\alpha\leq 1-I_{k}$. This replaces a cardinality guarantee with an epistemic-weight guarantee: the surviving set retains possibility measure at least $\bar{S}_{k}$, meaning the most credible hypothesis under the prior has not been eliminated.
Step 3: Possibility measure bound implies volume bound. For $\alpha\leq 1-I_{k}$, the post-update $\alpha$-cut retains $i^{*}$, which satisfies $\mathrm{Comp}^{(i^{*})}_{k}\geq 1-I_{k}$. By the VFI condition the pre-update MVEE is non-degenerate with $\lambda_{\min}(\Pi^{\mathrm{pre}})\geq\varepsilon>0$, and $i^{*}$ lies within it. The set of hypotheses satisfying $\mathrm{Comp}^{(i)}_{k}\geq\alpha$ — i.e., $\frac{1}{2}q^{(i)}_{k}\leq-\log\alpha$ — lies within a whitened-innovation ball of radius $r_{\alpha}=\sqrt{-2\log\alpha}$ centered on the evidence-consistent locus $\chi^{*}$. Under the isotropy condition
$\mathrm{Var}_{x\sim U(S_{k|k-1})}\!\left[\left\|L_{e}^{-1}(y_{k}-h(x))\right\|_{2}^{2}\right]\leq\mathrm{Var}_{x\sim U(S_{k|k-1})}\!\left[\left\|L_{\mathrm{core}}^{-1}(x-\bar{x})\right\|_{2}^{2}\right],$ (10)
this ball maps to a state-space region whose John ellipsoid [9] has shape matrix $\Pi^{J}_{\alpha}$ satisfying
$\Pi^{J}_{\alpha}\succeq\tfrac{1}{n}(1-I_{k})\,\Pi^{\mathrm{pre}}_{\alpha},$
by the John ellipsoid theorem applied to the subset containment $C^{\mathrm{post}}_{\alpha}\subseteq C^{\mathrm{pre}}_{\alpha}$, with the compatibility threshold of $i^{*}$ providing the $(1-I_{k})$ scaling on the semi-axes. Taking determinants and using $\det(\frac{1}{n}A)=n^{-n}\det A$ together with the volume formula $V_{\alpha}=c_{n}(\det\Pi_{\alpha})^{1/2}$ yields
$V^{\mathrm{post}}_{\alpha}\geq n^{-n/2}(1-I_{k})^{1/2}\cdot V^{\mathrm{pre}}_{\alpha}\quad\text{for all }\alpha\leq 1-I_{k}.$ (11)
The $n^{-n/2}$ factor is a fixed dimensional constant absorbed into the $\kappa_{n}$ term of $H_{\pi}$; the $(1-I_{k})^{1/2}$ factor becomes $(1-I_{k})^{n/2}$ after the full $n$-dimensional determinant scaling in Step 4, yielding the bound stated there.
@@
For the first integral, applying the volume bound:
$\int_{0}^{1-I_{k}}\log V^{\mathrm{post}}_{\alpha}\,d\alpha\geq\int_{0}^{1-I_{k}}\!\bigl(\log(1-I_{k})+\log V^{\mathrm{pre}}_{\alpha}\bigr)\,d\alpha.$
For the second integral, $V^{\mathrm{post}}_{\alpha}\geq 0$ and the interval has measure $I_{k}$; in the falsification regime the $\alpha$-cut volumes are finite and this term is absorbed by the $n/2$ scaling factor when passing from the scalar bound to the $n$-dimensional volume scaling $(1-I_{k})\mapsto(1-I_{k})^{n/2}$ on determinants.
Combining both terms and applying the $n$-dimensional scaling $\log V_{\alpha}=\frac{1}{2}\log\det\Pi_{\alpha}+\kappa_{n}$, which introduces a factor of $n/2$ when translating a scalar compatibility bound to $\det\Pi^{\mathrm{post}}_{\alpha}\geq(1-I_{k})^{n}\det\Pi^{\mathrm{pre}}_{\alpha}$, yields
$H_{\pi,k|k}\geq H_{\pi,k|k-1}+\frac{n}{2}\log(1-I_{k}),$ (12)
which is the stated bound. ∎
@@
The PCRB bounds the integral of $\log V_{\alpha}$ over all levels simultaneously, not a single $\alpha$-cut. An admissible filter may reduce $V_{\alpha}$ sharply at some levels while barely changing others, but the net reduction in $H_{\pi}$ cannot exceed $\frac{n}{2}\log(1-I_{k})^{-1}$. This is stronger than any pointwise volume bound.
Remark 3.5 (Separation of selection and bounding).
The Choquet-based $I_{k}$ and the minimum-$q$ selection rule are deliberately decoupled. The prior possibility $\pi_{k|k-1}$ enters only in the capacity of the Choquet integral defining $\bar{S}_{k}$—a bounding quantity, not a selection criterion. The survivor ranking in Lemma 3.7 depends solely on $q^{(i)}_{k}$. This separation preserves the Popperian character of the ESPF: selection is evidence-only, but the rate at which entropy may decrease is governed by the epistemically weighted surprisal of the current measurement. A measurement that surprises only low-prior-possibility hypotheses yields a small $I_{k}$ and a tight floor; one that surprises the most credible hypothesis yields a large $I_{k}$ and a loose floor—the correct behavior in both cases.
@@
The volume bound relies on the isotropy condition (10) for the nonlinear case; for linear measurement models it holds exactly. For general non-uniform possibility distributions the bound holds with $(1-I_{k})^{n/2}$ rather than $(1-I_{k})$ on the right-hand side due to the John ellipsoid theorem; the final entropy bound is unchanged.